Water vapour of mass 50g is pass through water of 5kg at 22 degree Celsius. Calculate the final temperature.
1
Expert's answer
2018-10-08T15:24:08-0400
A heat balance equation dives
Q_vapor=Q_water
mL+cm(100℃-t_final )=cM(t_final-t_initial)
So
t_final=(mL+cm×100+cMt_initial)/(c(m+M))
=(0.05×2.26×〖10〗^6+4200×0.05×100+4200×5×22)/(4200(0.05+5))=28 ℃